With what minimum acceleration can a fireman slide down a rope whose breaking strength is 2/5th of his weight?
1)1g
2).4g
3).6g
4).8g

Grade:12th pass

1 Answers

Eshan
5 years ago
Dear student,

When a person is sliding down,

ma=mg-T
\implies T=mg-ma\leq \dfrac{2}{5}mg
\implies a\geq\dfrac{3}{5}g=0.6g